A sample of helium gas behaves as an ideal gas as energy is added by heat at constant pressure from 273K t 393K. If 20.0J of work is done by the gas during this process, what is the mass of the helium present?


1
Expert's answer
2022-04-19T10:00:35-0400

The work done by gas during constant pressure process

"W=p\Delta V=\frac{m}{M}R\Delta T"

Hence, the mass of gas

"m=\frac{WM}{R\Delta T}"

"m=\frac{20.0*0.004}{8.31*(393-273)}=8.0*10^{-5}\:\rm kg"